# On $\operatorname{CAT}(\kappa)$ surfaces

### v2

@@ Abstract.
We study the properties of $\operatorname{CAT}(\kappa)$ surfaces: length metric spaces homeomorphic to a surface having curvature bounded above in the sense of satisfying the $\operatorname{CAT}(\kappa)$ condition locally. The main facts about $\operatorname{CAT}(\kappa)$ surfaces seem to be largely a part of mathematical folklore, and this paper is intended to rectify the situation. We provide three distinct proofs that $\operatorname{CAT}(\kappa)$ surfaces have bounded (integral) curvature. This fact allows one to apply the established theory of surfaces of bounded curvature to derive further properties of $\operatorname{CAT}(\kappa)$ surfaces. We also show that $\operatorname{CAT}(\kappa)$ surfaces can be approximated by smooth Riemannian surfaces of Gaussian curvature at most $\kappa$. We do this by giving explicit formulas for smoothing the vertices of model polyhedral surfaces.
@@ 1. Introduction
In the two-dimensional setting, there is a third class of spaces that were extensively studied by Alexandrov and the Leningrad school of geometry associated with him, namely surfaces of bounded (integral) curvature. Roughly speaking, this is the largest class of surfaces for which curvature can be meaningfully defined as a signed Radon measure. Its precise definition involves the notion of (angular) excess of a geodesic triangle $T$, denoted by $\delta(T)$. A surface of bounded curvature is one for which, in every compactly contained neighborhood, there is a uniform upper bound on the sum of excesses of an arbitrary finite collection of non-overlapping simple triangles. See Section 2 and Section 3 for complete definitions. There is a rich and well-developed theory of surfaces of bounded curvature developed in the monographs by Alexandrov–Zalgaller [1] and Reshetnyak [38] covering methods such as polyhedral approximation and conformal parametrizations; see also the surveys [19, 41].
@@ 1.1. The $\operatorname{CAT}(\kappa)$ condition and bounded curvature
This fact is perhaps not as well known as it could be. For example, the recent papers [8] and [14] prove results directly for $\operatorname{CAT}(\kappa)$ surfaces that are similar to more general results for bounded curvature surfaces, without explicitly recognizing the connection between the two classes of surfaces. 1.1 is trivial if $\kappa\leq 0$, since in this case every triangle has non-positive excess. In the case where $\kappa>0$ its validity is a priori not clear. We note that the analogue of 1.1 for surfaces with curvature bounded below is also true. A proof has been given by Richard in [39].
1.1 was already observed by Alexandrov himself in [3, Section 1.8] in his seminal work on $\operatorname{CAT}(\kappa)$ spaces, justified by an explanation to the effect that one can develop a parallel theory to that of bounded curvature surfaces for surfaces with curvature bounded above by $\kappa$. Alexandrov attributes the result to Zalgaller; however, an elaboration of this approach does not exist in the literature to our knowledge, and in any case seems to be more work than necessary. More recently, a separate proof of 1.1 was sketched by Machigashira–Ohtsuka in [30]. In fact, Lemma 5.3 in [30] states, in more generality, the following basic inequality relating excess and curvature, from which 1.1 follows easily.
@@
Let $T$ be a triangle in a $\operatorname{CAT}(\kappa)$ surface. Then $\delta(T)\leq\kappa|T|$.
Here, $|T|$ denotes the area of the interior of $T$, which we define as the Hausdorff $2$-measure. According to [30], 1.2 can be established in a similar manner to an analogous inequality for surfaces with curvature bounded below, which was done in earlier work of Machigashira [29, Theorem 2.0]. However, there are various qualitative differences between $\operatorname{CAT}(\kappa)$ spaces and spaces of curvature bounded below, and thus we feel that it is not straightforward to adapt the proof in [29] to the $\operatorname{CAT}(\kappa)$ case.
In light of the state of the literature, Fillastre asked in [17] for complete details of a proof. In this paper, we answer the request by giving three distinct proofs of 1.1, each of which has their own points of interest. These proofs are based on a recent theorem of the third-named author with Creutz [16] on decomposing an arbitrary surface with a length metric into non-overlapping simple convex triangles, which we refer to as a triangulation of the the surface. Note however that the second proof does not require the full strength of the triangulation theorem in [16], but rather the simpler fact that any point has a neighborhood composed of finitely many potentially overlapping triangles (Lemma 5.2 in [16]). The first proof is in our opinion the most comprehensive and intuitive; the main tool we use is a special type of triangulation that can be obtained by inductively subdividing a base triangulation, which we call a vertex-edge triangulation. The second is a more direct geometric proof, intended to establish 1.1 as efficiently as possible using two general facts, namely, Reshetnyak’s majorization theorem and the Besicovitch inequality. This proof is based on proving a non-sharp version of the inequality in 1.2. In the third proof, we verify instead an equivalent definition of being a limit of surfaces of uniformly bounded curvature. This proof is shorter than the others but instead verifies an alternative condition. We attribute the basic idea underlying the third proof to discussions with François Fillastre [18] and a MathOverflow post by Anton Petrunin [36].
Later in the paper, we show how 1.2 itself follows as a consequence of 1.1 using the theory of surfaces of bounded curvature. One technical point relates to the definition of area. As stated above, we take the Hausdorff $2$-measure as our definition. The original work of Alexandrov (cf. [3]) uses a different definition based on approximation by Euclidean polyhedral surfaces, which we call the Alexandrov area. We verify that the two notions of area are equivalent for $\operatorname{CAT}(\kappa)$ surfaces, or more generally for surfaces of bounded curvature without cusp points.
@@ 1.2. Smooth approximation
While one naturally expects 1.4 to be true, we have not found any complete statement in the literature. From the theory of surfaces of bounded curvature comes the fact that any $\operatorname{CAT}(\kappa)$ surface can be approximated by smooth surfaces with locally uniformly bounded integral curvature, though without any pointwise control on the curvature. See Chapter III of [1] and Section 6.2 of [38]. In [12, Theorem 0.11], Burago–Buyalo show that any $\operatorname{CAT}(\kappa)$ surface can be approximated by piecewise smooth $\operatorname{CAT}(\kappa)$ surfaces (more generally, their result covers any 2-dimensional locally $\operatorname{CAT}(\kappa)$ polyhedron). Recently, the case of 1.4 where $\kappa<0$ has been covered by a lemma of Labeni [24]. Labeni’s argument is based on the approximating cones with vertices of negative curvature in anti-de Sitter space by smooth convex surfaces.
To prove 1.4, the main task is to show that any vertex of a $\operatorname{CAT}(\kappa)$ polyhedral surface can be smoothed while retaining the $\operatorname{CAT}(\kappa)$ condition. We found that this can be accomplished via an explicit and relatively simple formula that interpolates between a metric of constant positive curvature away from the vertex with a metric of constant negative curvature around the vertex.
The analogue of 1.4 for surfaces of curvature bounded below has been addressed in full generality by Itoh–Rouyer–Vîlcu [21]; see also [37, 39] and the discussion in [22, 31] concerning the curvature bounded below case. Our approach also works for smoothing surfaces of curvature bounded below by $\kappa$, with the advantage of being more explicit than the approach in [21]; see 7.2. Thus we provide a concrete and uniform approach to the general problem of smoothing vertices of polyhedra.
@@ 1.3. The bigger picture
We now elaborate more on the consequences of 1.1 and place our work within a larger context. While the $\operatorname{CAT}(\kappa)$ condition provides an effective axiomatic condition for studying metric surfaces, there is a variety of increasing general conditions that have also been studied in the literature. Indeed, the $\operatorname{CAT}(\kappa)$ condition is fairly restrictive; roughly speaking, a $\operatorname{CAT}(\kappa)$ surface is one that is the limit of spherical polyhedral surfaces of uniformly bounded area with vertices of negative curvature. We state the following omnibus theorem, which is a compilation of easy or known facts, along with 1.1. For simplicity, we assume that the metric surface $X$ is closed.
@@
The implication (A) $\Longrightarrow$ (B) is classical; see the appendix to Chapter II.1 of [9]. The implication (B) $\Longrightarrow$ (C) is a restatement of 1.1 together with the simple observation that a $\operatorname{CAT}(\kappa)$ surface does not have any so-called cusp points, i.e., points having total curvature $2\pi$. Next, by a theorem of Reshetnyak [38, Theorem 9.10], the metric tangent space of every non-cusp point of a surface of bounded curvature is a Euclidean cone over a circle. That is, for each point $x$ on the surface and all $\varepsilon>0$ one can find $\delta>0$ such that the ball $B(x,r)$ is $(1+\varepsilon)$-bi-Lipschitz equivalent to a neighborhood of the above Euclidean cone. See Lemma 6 in [11] for a proof of Reshetnyak’s theorem. See also Bonk–Lang [6] for a strong quantitative version of this property. The implication (C) $\Longrightarrow$ (D) is shown by Burago in [11] using this local bi-Lipschitz equivalence. The implications (D) $\Longrightarrow$ (E) and (F) $\Longrightarrow$ (G) are elementary to verify from the definitions, while the implication (E) $\Longrightarrow$ (F) is due to Lytchak–Wenger [28, Corollary 5.5].
@@
One aspect of the theory concerns the topic of uniformization. The classical uniformization theorem states that any smooth Riemannian $2$-manifold is conformally equivalent to a surface of constant curvature. A major theorem of Reshetnyak concerns conformal parametrizations of bounded curvature surfaces: for any surface of bounded curvature, its metric can be represented locally by a length element of the form $ds=e^{\lambda}ds_{h}$, where $\lambda$ is the difference of two subharmonic functions and $ds_{h}$ is the length element of a background smooth Riemannian metric. The global version of this theorem is due to Huber [20]. There are similar uniformization theorems for the more general classes of surfaces in 1.5, including the Bonk–Kleiner theorem for Ahlfors $2$-regular, linearly locally contractible spheres [5]. A very general uniformization theorem applicable to any surface of locally finite Hausdorff $2$-measure, yielding a so-called weakly quasiconformal parametrization from a constant curvature surface, has been proved in [34, 35].
We refer the reader to the following references for various non-trivial examples of surfaces illustrating the difference between the classes of surfaces in 1.5: Laakso [23] for a surface satisfying (E) but not (D); Theorem 1.1 in [15] for a surface satisfying (F) but not (E); Theorem 1.2 in [40] for a surface satisfying (G) but not (F). In addition, we observe that $\mathbb{R}^{2}$ equipped with any normed metric not arising from an inner product satisfies (D) but not (C). Moreover, any polyhedral surface with cone points of positive curvature satisfies (C) but not (B).
@@ 1.4. Outline of the paper
The paper is organized as follows. We collect preliminaries related to metric geometry in Section 2 and related to $\operatorname{CAT}(\kappa)$ spaces in Section 3. The material on $\operatorname{CAT}(\kappa)$ spaces is probably well-known to experts but includes some propositions that we could not find in the literature. Section 4 describes our main tool of vertex-edge triangulations in preparation for the first proof of 1.1. The three proofs of 1.1 are given in Section 5. The other results are proved in the following sections. First, we verify 1.3 on definitions of area and 1.2 relating the excess of a triangle to its area in Section 6. Finally, we prove 1.4 on smooth approximation in Section 7.
@@ 2.1. Basics of metric geometry
A useful fact about Hausdorff measure is the Besicovitch inequality. We state a two-dimensional version for metric spaces. See Theorem 2.1 in [34] for a proof.
Theorem 2.1 (Besicovitch inequality).
Let $X$ be a metric space homeomorphic to a closed disk with topological boundary $\partial X$. Suppose there is a closed Jordan domain $\Omega\subset\mathbb{R}^{2}$ and a $L$-Lipschitz bijective map from $\partial X$ to $\partial\Omega$. Then
$|X|\geq\frac{\pi|\Omega|}{4L^{2}}.$
@@ 2.2. Polygons and triangulations
We say that a simple polygon $P$ is a Jordan polygon if it is contained in a neighborhood $U$ homeomorphic to an open disk and the diameter of $P$ is at most $d(P,\partial U)$. The last requirement guarantees that the length metric restricted to $P$ is the same in $U$ as in the original space. By the Jordan curve theorem, any simple polygon $P$ is the boundary of a closed Jordan domain contained in the ambient neighborhood $U$. The interior of this domain is called the interior of the polygon, denoted by $P^{\circ}$. Note that the previous fact depends on our assumption of an ambient disk neighborhood; otherwise, a simple polygon may not bound a unique disk, or any disk at all. For example, a great circle on the $2$-sphere can be viewed as a triangle, and either hemisphere may be designated as the interior.
@@
Let $X$ be a length surface and $\mathcal{U}$ a cover of $X$ by topological open disks. For all $\varepsilon>0$ there exists a triangulation $\mathcal{T}$ of $X$ with mesh at most $\varepsilon$ and that each element $T\in\mathcal{T}$ is contained in a set $U\in\mathcal{U}$ satisfying $d(T,\partial U)\geq\varepsilon$. Moreover, if $\mathcal{E}$ is the edge graph of $\mathcal{T}$ with the induced length metric, then the inclusion of $\mathcal{E}$ into $X$ is an $\varepsilon$-isometry.
@@
The statement concerning the open cover $\mathcal{U}$ is not part of the triangulation theorem as stated in [16], although it follows from Remark 5.5 in [16]. Also, the triangulation theorem in [16] gives the stronger property of boundary convexity rather than just convexity, though this is not needed for our purposes. The final conclusion regarding the edge graph can be found in Proposition 5.2 in [34].

### v3

@@ Abstract.
We study the properties of $\operatorname{CAT}(\kappa)$ surfaces: length metric spaces homeomorphic to a surface having curvature bounded above in the sense of satisfying the $\operatorname{CAT}(\kappa)$ condition locally. The main facts about $\operatorname{CAT}(\kappa)$ surfaces seem to be largely a part of mathematical folklore, and this paper is intended to rectify the situation. We provide a complete proof that $\operatorname{CAT}(\kappa)$ surfaces have bounded (integral) curvature. This fact allows one to apply the established theory of surfaces of bounded curvature to derive further properties of $\operatorname{CAT}(\kappa)$ surfaces. We also show that $\operatorname{CAT}(\kappa)$ surfaces can be approximated by smooth Riemannian surfaces of Gaussian curvature at most $\kappa$. We do this by giving explicit formulas for smoothing the vertices of model polyhedral surfaces.
@@ 1. Introduction
In the two-dimensional setting, there is a third class of spaces that were extensively studied by Alexandrov and the Leningrad school of geometry associated with him, namely surfaces of bounded (integral) curvature. Roughly speaking, this is the largest class of surfaces for which curvature can be meaningfully defined as a signed Radon measure. Its precise definition involves the notion of (angular) excess of a geodesic triangle $T$, denoted by $\delta(T)$. A surface of bounded curvature is one for which, in every compactly contained neighborhood, there is a uniform upper bound on the sum of excesses of an arbitrary finite collection of non-overlapping simple triangles. See Section˜2 and Section˜3 for complete definitions. There is a rich and well-developed theory of surfaces of bounded curvature; see the monographs by Alexandrov–Zalgaller [1] and Reshetnyak [36] covering methods such as polyhedral approximation and conformal parametrizations, as well as the recent surveys [17, 39].
@@ 1.1. The $\operatorname{CAT}(\kappa)$ condition and bounded curvature
This fact is perhaps not as well known as it could be. For example, some of the theorems in the recent papers [6] and [13] on $\operatorname{CAT}(\kappa)$ surfaces are essentially special cases of more general results for bounded curvature surfaces, and the connection between these two classes of surfaces is not made. ˜1.1 is trivial if $\kappa\leq 0$, since in this case every triangle has non-positive excess. In the case where $\kappa>0$ its validity is a priori not clear. We note that the analogue of ˜1.1 for surfaces with curvature bounded below is also true. A proof has been given by Richard in [37].
˜1.1 was already observed by Alexandrov himself in [2, Section 1.8] in his seminal work on $\operatorname{CAT}(\kappa)$ spaces, justified by an explanation to the effect that one can develop a parallel theory to that of bounded curvature surfaces for surfaces with curvature bounded above by $\kappa$. Alexandrov attributes the result to Zalgaller; however, an elaboration of this approach does not exist in the literature to our knowledge, and in any case seems to be more work than necessary. More recently, a separate proof of ˜1.1 was sketched by Machigashira–Ohtsuka in [28]. In fact, Lemma 5.3 in [28] states, in more generality, the following basic inequality relating excess and curvature, from which ˜1.1 follows easily.
@@
Let $T$ be a triangle in a $\operatorname{CAT}(\kappa)$ surface that is the boundary of a closed disk. Then $\delta(T)\leq\kappa|T|$.
Here, $|T|$ denotes the area of the interior of $T$, which we define as the Hausdorff $2$-measure of this set. According to [28], ˜1.2 can be established in a similar manner to an analogous inequality for surfaces with curvature bounded below, which was done in earlier work of Machigashira [27, Theorem 2.0]. However, there are various qualitative differences between $\operatorname{CAT}(\kappa)$ spaces and spaces of curvature bounded below, and thus we feel that it is not straightforward to adapt the proof in [27] to the $\operatorname{CAT}(\kappa)$ case.
In light of the state of the literature, Fillastre asked in [16] for complete details of a proof. In this paper, we answer the request by giving a complete proof of ˜1.1. Our approach is based on a recent theorem of the third-named author with Creutz [15] on decomposing an arbitrary surface with a length metric into non-overlapping simple convex triangles, which we refer to as a triangulation of the the surface. See ˜2.4 below. The main tool we use is a special type of triangulation that can be obtained by inductively subdividing a base triangulation, which we call a vertex-edge triangulation.
We remark that $\operatorname{CAT}(\kappa)$ surfaces can be equivalently defined as limits of smooth or polyhedral surfaces of locally uniformly bounded curvature. As an alternative proof of ˜1.1, it is possible to use the triangulation theorem in [15] to verify this equivalent condition. However, the proof we give in this paper seems to us the most intuitive and comprehensive.
Later in the paper, we show how ˜1.2 itself follows readily as a consequence of ˜1.1 using the theory of surfaces of bounded curvature. One technical point relates to the definition of area. As stated above, we take the Hausdorff $2$-measure as our definition. The original work of Alexandrov (cf. [2]) uses a different definition based on approximation by Euclidean polyhedral surfaces, which we call the Alexandrov area. We verify that the two notions of area are equivalent for $\operatorname{CAT}(\kappa)$ surfaces, or more generally for surfaces of bounded curvature without cusp points:
@@ 1.2. Smooth approximation
While one naturally expects ˜1.4 to be true, we have not found any complete statement in the literature. From the theory of surfaces of bounded curvature comes the fact that any $\operatorname{CAT}(\kappa)$ surface can be approximated by smooth surfaces with locally uniformly bounded integral curvature, though without any pointwise control on the curvature. See Chapter III of [1] and Section 6.2 of [36]. In [11, Theorem 0.11], Burago–Buyalo show that any $\operatorname{CAT}(\kappa)$ surface can be approximated by piecewise smooth $\operatorname{CAT}(\kappa)$ surfaces (more generally, their result covers any 2-dimensional locally $\operatorname{CAT}(\kappa)$ polyhedron). Recently, the case of ˜1.4 where $\kappa<0$ has been covered by a lemma of Labeni [22]. Labeni’s argument is based on the approximating cones with vertices of negative curvature in anti-de Sitter space by smooth convex surfaces.
To prove ˜1.4, the main task is to show that any vertex of a $\operatorname{CAT}(\kappa)$ polyhedral surface can be smoothed while retaining the $\operatorname{CAT}(\kappa)$ condition. We found that this can be accomplished via an explicit and relatively simple formula that interpolates between a metric of constant positive curvature away from the vertex with a metric of constant negative curvature around the vertex.
The analogue of ˜1.4 for surfaces of curvature bounded below has been addressed in full generality by Itoh–Rouyer–Vîlcu [19]; see also [34, 37] and the discussion in [20, 29] concerning the curvature bounded below case. Our approach also works for smoothing surfaces of curvature bounded below by $\kappa$, with the advantage of being more explicit than the approach in [19]; see ˜7.2. Thus we provide a concrete and uniform approach to the general problem of smoothing vertices of polyhedra.
@@ 1.3. The bigger picture
We now elaborate more on the consequences of ˜1.1 and place our work within a larger context. While the $\operatorname{CAT}(\kappa)$ condition provides an effective axiomatic condition for studying metric surfaces, there is a variety of increasingly general conditions that have also been studied in the literature. Indeed, the $\operatorname{CAT}(\kappa)$ condition is fairly restrictive; roughly speaking, a $\operatorname{CAT}(\kappa)$ surface is one that is the limit of spherical polyhedral surfaces of uniformly bounded area with vertices of negative curvature. We state the following omnibus theorem, which is a compilation of easy or known facts, along with ˜1.1. For simplicity, we assume that the metric surface $X$ is closed.
@@
The implication (A) $\Longrightarrow$ (B) is classical; see the appendix to Chapter II.1 of [7]. The implication (B) $\Longrightarrow$ (C) is a restatement of ˜1.1 together with the simple observation that a $\operatorname{CAT}(\kappa)$ surface does not have any so-called cusp points, i.e., points having total curvature $2\pi$. Next, the implication (C) $\Longrightarrow$ (D) follows from a theorem of Reshetnyak that surfaces of bounded curvature without cusp points can be approximated in the bi-Lipschitz sense by polyhedral surfaces; see Theorems 9.10 and 9.11 in [36] and Lemma 6 in [10]. See also Bonk–Lang [4] for a strong quantitative version of this property. The implications (D) $\Longrightarrow$ (E) and (F) $\Longrightarrow$ (G) are elementary to verify from the definitions, while the implication (E) $\Longrightarrow$ (F) is due to Lytchak–Wenger [26, Corollary 5.5].
@@
One aspect of the theory concerns the topic of uniformization. The classical uniformization theorem states that any smooth Riemannian $2$-manifold is conformally equivalent to a surface of constant curvature. A major theorem of Reshetnyak concerns conformal parametrizations of bounded curvature surfaces: for any surface of bounded curvature, its metric can be represented locally by a length element of the form $ds=e^{\lambda}ds_{h}$, where $\lambda$ is the difference of two subharmonic functions and $ds_{h}$ is the length element of a background smooth Riemannian metric [35]. The global version of this theorem is due to Huber [18]. There are similar uniformization theorems for the more general classes of surfaces in ˜1.5, including the Bonk–Kleiner theorem for Ahlfors $2$-regular, linearly locally contractible spheres [3]. A very general uniformization theorem applicable to any surface of locally finite Hausdorff $2$-measure, yielding a so-called weakly quasiconformal parametrization from a constant curvature surface, has been proved in [32, 33].
We refer the reader to the following references for various non-trivial examples of surfaces illustrating the difference between the classes of surfaces in ˜1.5: Laakso [21] for a surface satisfying (E) but not (D); Theorem 1.1 in [14] for a surface satisfying (F) but not (E); Theorem 1.2 in [38] for a surface satisfying (G) but not (F). In addition, we observe that $\mathbb{R}^{2}$ equipped with any normed metric not arising from an inner product satisfies (D) but not (C). Moreover, any polyhedral surface with cone points of positive curvature satisfies (C) but not (B).
@@ 1.4. Outline of the paper
The paper is organized as follows. We collect preliminaries related to metric geometry in Section˜2 and related to $\operatorname{CAT}(\kappa)$ spaces in Section˜3. The material on $\operatorname{CAT}(\kappa)$ spaces is probably well-known to experts but includes some propositions that we could not find in the literature. Section˜4 describes our main tool of vertex-edge triangulations in preparation for the proof of ˜1.1. The proof of ˜1.1 is given in Section˜5. The other results are proved in the following sections. First, we verify ˜1.3 on definitions of area and ˜1.2 relating the excess of a triangle to its area in Section˜6. Finally, we prove ˜1.4 on smooth approximation in Section˜7.
@@ 2.2. Polygons and triangulations
We say that a simple polygon $P$ is a Jordan polygon if it is contained in a neighborhood $U$ homeomorphic to an open disk. By the Jordan curve theorem, any simple polygon $P$ is the boundary of a closed Jordan domain contained in the ambient neighborhood $U$. The interior of this domain is called the interior of the polygon, denoted by $P^{\circ}$. Note that the previous fact depends on our assumption of an ambient disk neighborhood; otherwise, a simple polygon may not bound a unique disk, or any disk at all. For example, a great circle on the $2$-sphere can be viewed as a triangle, and either hemisphere may be designated as the interior.
@@
Let $X$ be a length surface and $\mathcal{U}$ a cover of $X$ by open sets. For all $\varepsilon>0$ there exists a triangulation $\mathcal{T}$ of $X$ with mesh at most $\varepsilon$ such that each element $T\in\mathcal{T}$ is contained in a set $U\in\mathcal{U}$ satisfying $d(T,\partial U)>\operatorname{diam}T$. Moreover, if $\mathcal{E}$ is the edge graph of $\mathcal{T}$ with the induced length metric, then the inclusion of $\mathcal{E}$ into $X$ is an $\varepsilon$-isometry.
The same result holds for surfaces with boundary provided the boundary is piecewise geodesic.
@@
The statements concerning the open cover $\mathcal{U}$ is not part of the triangulation theorem as stated in Theorem 1.2 of [15], although it follows from Remark 5.5 in [15]. Also, the triangulation theorem in [15] gives the stronger property of boundary convexity rather than just convexity, though this is not needed for our purposes. The final conclusion regarding the edge graph can be found in Proposition 5.2 in [32].